How much concrete should I order to cover 48' x 2' x 3"?
Answer by jim_thompson5910(35256)

3 inches = 3/12 = 0.25 feet

So the volume of the 48' by 2' by 0.25' slab is 48*2*0.25 = 24 cubic feet